OpenStudy (anonymous):

3-4(z+1)=3z-(2+z) solve for z PLZ HELP

OpenStudy (studygurl14):

3 - 4(z-4) = 3z - (2 + z) 3 - 4z + 16 = 3z -2 + z -4z + 19 = 4z - 2 You do the rest

OpenStudy (anonymous):

yes plz

OpenStudy (studygurl14):

yes plz what? You should be able to do the rest by your self. If not, then tell me what your having trouble with

OpenStudy (anonymous):

nothing lol

OpenStudy (studygurl14):

Okay then. are you okay doing the rest?

OpenStudy (anonymous):

yea thx

OpenStudy (studygurl14):

'welcome